OpenStudy (aaronandyson):

if 2y - 4x = 5 and 4x + ky + 3 = 0 are perpendicular to each other,find the value of k

OpenStudy (aaronandyson):

@campbell_st

OpenStudy (anonymous):

M1 X M2 = -1

OpenStudy (campbell_st):

ok... so the line 2y - 4x = 5 can you write it in slope intercept form y = mx + b..?

OpenStudy (aaronandyson):

M1 = 2??

OpenStudy (campbell_st):

well ok... so what is the negative reciprocal of 2...?

OpenStudy (aaronandyson):

-2

OpenStudy (anonymous):

reciprocal

OpenStudy (campbell_st):

well thats the negative... reciprocal means to turn it upside down so 2/1 becomes 1/2 so that means the slope of the perpendicular line is -1/2 now take the like 4x + ky + 3 = 0 can you write this in slope intercept form..?

OpenStudy (campbell_st):

as a start ky = -4x - 3 what do you think the slope of this line is..?

OpenStudy (aaronandyson):

-4/k

OpenStudy (anonymous):

for the line to be perpendicular m1 x m2 = -1

OpenStudy (aaronandyson):

k= 8